A non-dividend paying stock X is trading at $10x. A put option for 3 years maturity on the non-dividend paying stock X at a strike price of $8z costs $2t. An investor decides to enter in a protective put position for this stock for 3 years. Calculate undiscounted profit or loss of the investor at the end of the maturity if the terminal spot price of stock X turns out to be $8t.

if you see x,y,z,t you can use it like that

x=5, y=3, z=6, t=5

Round to at least 6 decimals unless otherwise stated
Answer #1

Undiscounted profit=MAX(8z-8t,0)-2t+8t-10x=MAX(8*6-8*5,0)-2*5+8*5-10*5=-12.00

Loss of $12